OpenStudy (anonymous):

write each indicated quotient as a single complex number. 3-2i/2+i

OpenStudy (anonymous):

You got to multiply top and bottom with the complex conjugate of the denominator. Its like how you simplify surds.

OpenStudy (anonymous):

The complex conjugate of a complex number is simply the same complex number except that the imaginary part is multiplied by negative one.

OpenStudy (mathmath333):

use this idea \(\large\tt \begin{align} \color{black}{\dfrac{3-2i}{2+i}\\~\\ =\dfrac{3-2i}{2+i}\times\dfrac{2-i}{2-i}\\~\\ }\end{align}\)